Jai alai balls come in boxes of 8 and 15, so that 38 balls (one small box and two large boxes) can be purchased without having to break open a box, but 37 balls cannot. Find the maximum number of balls that cannot be bought without breaking boxes
Note: sorry didn’t know which category to put it in

Answer by greenestamps(13241)



Given two positive integers a and b that are relatively prime, the largest sum that cannot be made using multiples of a and b is the product of the two integers, minus the sum of the two integers: (ab)-(a+b).

For this problem, with the integers 8 and 15, the largest sum that cannot be made is (8*15)-(8+15) = 120-23 = 97.